### v2

# DivShift: Exploring Domain-Specific Distribution Shift in Volunteer-Collected Biodiversity Datasets

Abstract
Climate change is negatively impacting the world’s biodiversity. To build automated systems to monitor these negative biodiversity impacts, large-scale, volunteer-collected datasets like iNaturalist are built from community-identified, natural imagery. However, such volunteer-based data are opportunistic and lack a structured sampling strategy, resulting in geographic, temporal, observation quality, and socioeconomic, biases that stymie uptake of these models for downstream biodiversity monitoring tasks. Here we introduce DivShift North American West Coast (DivShift-NAWC), a curated dataset of almost 8 million iNaturalist plant images across the western coast of North America, for exploring the effects of these biases on deep learning model performance. We compare model performance across four known biases and observe that they indeed confound model performance. We suggest practical strategies for curating datasets to train deep learning models for monitoring climate change’s impacts on the world’s biodiversity.
**footnotetext: Denotes equal contribution to this work
1 Introduction
The world’s biodiversity is under threat from climate and land-use change [96, 90]. Biodiversity helps ecosystems combat climate change by improving carbon sequestration [137] and novel climate adaptation [105]. Therefore, monitoring the world’s biodiversity via automated tools [100] is critical to mitigate climate change’s effects on the natural world. Building machine learning tools for this automated monitoring requires large volumes of natural world imagery [53]. Participatory science applications—where users can upload and identify photos of species in their natural environments—have surged in popularity [37, 3, 39] (see Appdx. 5.1.1 for participatory science overview). These applications now provide sufficient finely-labeled imagery data to build large-scale biodiversity datasets [125, 60, 122, 107, 50, 55] and train deep learning models for automated biodiversity monitoring tasks such as species recognition [125, 50, 55, 60], species distribution modeling [122, 60, 107, 53], novel species identification [115], and visual question answering [125].
Figure 1: Biodiversity data biases include spatial bias, e.g. more observations in urban compared to wild areas; temporal bias, e.g. more observations on weekends compared to weekdays; observation quality bias, e.g. more diverse observations made by highly engaged observers than infrequent users; and sociopolitical bias, e.g. disparities between observation density due to political as opposed to ecological boundaries.
However, as these observations become easier for the public to make, sampling becomes more unstructured [2, 98, 52] and injects biases into these data [63, 37, 11, 39]. Therefore, volunteer-collected biodiversity datasets often do not fully reflect the world’s biodiversity [3, 27, 20, 70], presenting challenges for the general uptake of these data for biodiversity monitoring [66, 12, 34]. To help quantify the effects of these biases on model performance, here we introduce a new public biodiversity imagery dataset DivShift-North American West Coast (DivShift-NAWC), comprised of  8 million observations and 10,000 plant species across the North American west coast. DivShift-NAWC spans wide climatic, ecological, and sociopolitical gradients, enabling the targeted testing of downstream performance effects of spatial, temporal, observation quality, and sociopolitical biases present in volunteer-collected biodiversity datasets.
2 Related Works
Biases in volunteer-collected biodiversity datasets. Collection of large-scale volunteer datasets are subject to social and ecological filters [21, 63], which inject many types of bias into biodiversity datasets built from these collections [21, 38, 64] (see Appdx. 5.1.2 for further definition). In this work, we focus on four kinds of bias common to volunteer-collected biodiversity datasets: spatial, temporal, observation quality, and sociopolitical (Fig. 1). Spatial bias includes observer preferences to sampling easy-to-access greenspaces in urban areas [56, 83, 3, 38, 39, 21] (see Appdx. 5.1.3). Temporal bias [3, 39, 119, 106, 30, 61] includes a skew towards more observations on weekends when observers are free from work [38, 29, 27] (see Appdx. 5.1.3). Observation quality bias [39, 124, 85, 10] manifests as a small but dedicated group of users that tend to observe more species in more diverse habitats [38, 104] (see Appdx. 5.1.3). Lastly, sociopolitical bias in who has access to the resources, time, and areas to collect biodiversity observations [9, 41, 80, 34, 24, 21, 16, 26, 111, 93] includes a skew towards whiter, wealthier, and older observers [79, 97] (see Appdx. 5.1.3).
Large-scale natural world imagery datasets. Large-scale natural world imagery datasets for training computer vision models for biodiversity monitoring tasks span a variety of modalities, including handheld phone images [125, 60, 107, 50, 55, 126], high-quality archival and herbaria images [115, 33], long-distance camera imagery [131, 71], terrestrial camera traps [6, 118], ocean sonar cameras [69], google street view imagery [7, 73] and remote sensing imagery [25, 53, 122, 60, 132]. With these datasets, efforts to minimize the domain-specific bias of biodiversity data range from using context to differentially select training examples [94], mixing high-quality, expert-curated images in with volunteer-collected images [115], evenly sampling images by class [125], and spatial stratification [60, 25, 53]. However, none of these datasets nor techniques explicitly address how each type of bias present in these datasets affect downstream model performance.
3 DivShift Framework and DivShift-NAWC Dataset
To explore the downstream effects of bias in volunteer-collected biodiversity datasets, we first propose the bioDiversity Shift (DivShift) framework, then introduce the DivShift North American West Coast (DivShift-NAWC) dataset.
DivShift Framework: The bioDiversity Shift (DivShift) framework casts biases present in volunteer-collected biodiversity data as distribution shifts. Specifically, the dataset of individuals actually uploaded by observers, $D$, is the result of the biased sampling process $\stackrel{{\scriptstyle S_{a}}}{{\sim}}$ from the true distribution of biodiversity, $J$, leading to strong skews in $D$ across space, time, taxon, observer, and sociopolitical boundaries. To explore these differences, we partition $D$ into $P_{A}$ and $P_{B}$ by some known bias in biodiversity data, (e.g. for spatial bias $P_{A}$ being observations from modified cities and $P_{B}$ being observations from undisturbed wilderness). We then further sub-partition observations in $P_{A}$ and $P_{B}$ randomly into 80% train $P_{Atrain}$ and 20% $P_{Atest}$ test as normally done in machine learning datasets [35, 125, 7], then measure the underlying effect of this distribution shift via the Jensen-Shannon Distance (JSD) between $P_{Atrain}$ and $P_{Btest}$. Finally, we compare the JSD to the change in performance between deep learning models trained on $P_{Atrain}$ and tested on $P_{Atest}$ to the same models tested on $P_{Btest}$. If a model’s performance decreases when tested on $P_{Btest}$, we consider that to be a negative bias, and a positive bias if it increases. Moreso, if the change in a model’s performance is less than the underlying JSD between the partitions $P_{Atrain}$ and $P_{Btest}$, it is considered a weak bias, while if it is greater, it is considered a strong bias.
DivShift North American West Coast Dataset: To test these distribution effects, we use the new DivShift - North American West Coast (DivShift-NAWC) dataset. DivShift-NAWC consists of 7.3 million iNaturalist images from the west coast of North America between 2019 and 2023, which spans three countries, eleven states, and eleven ecosystems (Table A1, Fig. 2). Biodiversity data by nature is long-tailed [43], and so like all other popular natural world imagery datasets [126, 125, 7, 60], DivShift-NAWC is heavily long-tailed.
We focus on spatial, temporal, observation quality, and sociopolitical bias (Fig. 1). For spatial bias, we use the Human Footprint Index (HFI) [87] where observations in wilderness are HFI <= 1 and modified are HFI >= 4 (Appdx. 5.3.2). For temporal bias we compare observations from the City Nature Challenge (CNC) [37, 61] to those not (Appdx. 5.3.3). For observation quality bias, compare engaged observers ($>$ 1,000 observations) to casual observers ($<$ 50 observations) [38] (Appdx. 5.3.4). Finally for sociopolitical bias, we compare two high-resource states ($>1$ mil. images, California-C) and British Columbia-BC) to two low-resource states ($<50,000$ images, Sonora-SO and Yukon-YT) [102] (Appdx. 5.3.5). We also report accuracies on baseline train / test partitions as commonly executed in other natural world imagery datasets [125, 25, 60, 35] (Appdx. 5.3.6).
Figure 2: Overview of DivShift-North American West Coast Dataset. a. Density plot of the dataset’s iNaturalist observations [89]. Observations are skewed to U.S. and coastal states. b. DivShift-NAWC spans a diverse set of habitats and ecosystems [92], c. along with climates [136]. d. DivShift-NAWC observations are concentrated in human-modified areas [87].
Model Training and Evaluation Metrics: To measure the underlying data partition distribution shifts, we filter each bias partition (e.g. modified cities vs. wilderness) to only species shared between the partitions, and calculated the Jensen-Shannon Distance (JSD) between the label distributions (Appdx. 5.5). For each bias partition, we split images into either train (80%) or test (20%) unless otherwise specified, and train a ResNet18 for 10 epochs on a given train partition (see Appdx. 5.6 for training details). For each model, we report Top1 accuracy per-observation (Top1-Obs), Top1 accuracy per-species (Top1-Spec), a new Top1 accuracy metric weighted by rarity (Top1-Wgt), and Top1 accuracy by land use category (Top1-LUC) (see Appdx. 5.7 for metric definitions).
4 Results and Discussion
For all bias partitions, surprisingly all biases are weak, with performance drop greater than the underlying JSD of the partition (Tables 1(a), 1(b)). For the spatial partition, there is a larger distribution shift from wild to modified observations than vice-versa (Table 1(b)). However, training on observations from less-disturbed habitats leads to worse performance than training in areas of high human activity (Table 1(a)). For the temporal partition, the distribution shift is largely symmetric (Table 1(b)), but training on City Nature leads universally worse performance on observations from outside the Challenge, while training with observations from outside the Challenge leads to better performance except when accounting for land use (Table 1(a)). For the observation quality partition we see that the distribution shift between both the engaged and casual observer partitions are surprisingly symmetric (Table 1(b)). However, the model trained using images from engaged observers showed increased accuracy across the board, while the model trained on casual observations saw a universal decrease (Table 1(a)). For the sociopolitical bias, the distribution shift is greater for CA to SO than BC to YT (Table 1(b)), yet surprisingly the decrease in performance is less pronounced for CA to SO than BC to YT, except when correcting for land use type (Table 1(a)). Across the four splits, model accuracies (Table A3) are well within the range of the baseline splits (Table A2), implying that the bias splits are not dramatically skewed compared to previous approaches.
Train-Test Diff. Top1-Obs Top1-Spec Top1-Wgt Top1-LUC
Spatial Bias (Human Footprint)
Wild-Modified diff -0.353 -0.154 -0.076 -0.376
Modified-Wild diff -0.113 0.068 0.322 -.111
Temporal Bias (City Nature Challenge)
CNC-Not CNC diff -0.179 -0.100 -0.065 -0.164
Not CNC-CNC diff 0.007 0.091 0.200 -0.009
Observer Quality Bias
Casual-Engaged diff -0.231 -0.125 -0.067 -0.192
Engaged-Casual diff 0.049 0.038 0.097 0.29
Sociopolitical Bias
CA-SO diff -0.248 0.014 0.294 -0.336
BC-YT diff -0.309 -0.095 0.069 -0.284
Train - Test JSD
Spatial Bias (Human Footprint)
Wild - Modified 0.643
Modified - Wild 0.618
Temporal Bias (City Nature)
CNC - Not CNC 0.307
Not CNC - CNC 0.296
Observation Quality Bias
Casual - Engaged 0.376
Engaged - Casual 0.372
Socio-Political Bias
CA - SO 0.571
BC - YT 0.518
Table 1: (1(a)) Performance differences for all partitions (absolute values in Table A3). Blue values indicate positive bias, red values indicate negative bias. Obs = Observation, Spec=Species, Wgt=Weighted, LUC=Land Use Category, CNC = City Nature Challenge, diff = difference. (1(b)) Jensen-Shannon differences between partitions. JSD = Jensen-Shannon Difference.
Climate Change Impact: Our findings on the DivShift-NAWC dataset suggest four recommendations for training deep learning models on voluminous but noisy volunteer-collected biodiversity datasets. First, observations from urban areas provide useful training signal even for wild areas for species found in both kinds of habitats. Second, using many noisy observations with a random sampling pattern is better than using fewer observations with a more structured sampling pattern. Third, using data from more engaged observers is generally better than less-engaged ones. Fourth, even though sociopolitical boundaries can have a significant effect on performance, training on diverse and extremely well-sampled high-resource regions like CA can still provide some predictive benefit for low-resource regions even if geographically far away.
Conclusion Here we introduce DivShift-NAWC, a new large-scale natural world imagery dataset designed to benchmark distribution shift effects on computer vision models for biodiversity monitoring. This framework and dataset enable the rigorous testing of problems known to the conservation biology community in a machine learning setting to help enable the building of more robust, accurate biodiversity monitoring tools for counteracting climate change’s effects on the world’s biodiversity.
Acknowledgments and Disclosure of Funding
We first and foremost thank all of the participants who volunteered and contributed observations on iNaturalist. We further thank Noah Goodman and Gabriel Poesia for their comments and discussion, and we further thank CoCoLab for donating compute for this project. We also thank the TomKat Center for Sustainable Energy, The Fulbright Brasil Commission, Carnegie Science, the Howard Hughes Medical Institute, and the University of California, Berkeley for funding support for this research. This research was also funded by the NSF Graduate Research Fellowship DGE-1656518 (L.G.), the TomKat Graduate Fellowship for Translational Research (L.G.), and the Krupp Internship Program for Stanford Students in Germany (E.S.). T.K. and S.S. acknowledge funding from the German Research Foundation (DFG) within the framework of BigPlantSens (Assessing the Synergies of Big Data and Deep Learning for the Remote Sensing of Plant Species; project no. 444524904) and PANOPS (Revealing Earth’s plant functional diversity with citizen science; project no. 504978936). This research was further funded by the Office of the Director of the National Institutes of Health’s Early Investigator Award (1DP5OD029506-01, M.E.-A.). Compute for this project was performed on the Calc cluster at Carnegie Science and the Stanford SC Compute Cluster.